# Smoothing Variances Across Time: Adaptive Stochastic Volatility

### v4

@@ 5 Empirical Study
We continue our analysis of the datasets introduced in Section 1 and estimate the log-variance using four methods: SV and MSSV2, which serve as established baselines, and ASV-DHS and ASV-DHS-N, our proposed models. For the ASV models, we set $k=1$ to reflect the presence of abrupt changes observed in Figure 1. In the S&P 500 dataset (Figures 4(a), 4(b), 4(c) and 4(d)), all models detect major volatility spikes in 2009 and 2020, corresponding to the global financial crisis and the COVID-19 pandemic. Moderate increases are also identified around 2001 (9/11), 2011 (European debt crisis), and 2018 (U.S. government shutdowns). ASV-DHS and ASV-DHS-N provide smoother and more stable estimates compared to SV and MSSV2, particularly during low-volatility periods.
For both the electricity (Figures 4(e), 4(f), 4(g) and 4(h)), and bike rental datasets (Figures 4(i), 4(j), 4(k) and 4(l)), all methods capture strong seasonal patterns in volatility, with elevated uncertainty coinciding with periods of peak activity. In the electricity dataset, this corresponds to high-demand months, while in the bike rental dataset, volatility increases during summer.
@@
Table 3: Summary statistics for each dataset and model: (1) the mean and (2) kurtosis of the estimated log-variance increments $|\Delta\hat{h}_{t}|$ and (3) the total number of flagged change points (CP). CP are computed as $\sum_{t=2}^{T}\mathds{1}\{|\Delta\hat{h}_{t}|>\delta\}$, where $\delta:=5\times\texttt{sd}(\Delta\hat{h}_{t})$.
Table 3 presents summary metrics of the estimated log-variance increments $|\Delta\hat{h}_{t}|$ across the three empirical datasets. These metrics, mean, kurtosis, and the number of flagged change points (CP), offer insight into the smoothness, tail behavior, and local adaptivity. ASV-DHS consistently produces the lowest mean, indicating highly smooth volatility trajectories. At the same time, it shows extremely high kurtosis and the highest number of flagged change points, suggesting heavy-tailed behavior in the increments and strong local adaptivity to sudden structural shifts. These properties highlight the model’s tendency to promote overall smoothness while remaining responsive to large isolated changes.
The smoothness of ASV-DHS-N adapts to the structure of the data. In the electricity dataset, where the volatility is characterized by extended periods of small changes punctuated by occasional abrupt shifts, the mean in Table 3 under ASV-DHS-N is similar to that of ASV-HS. In contrast, for the bike rental and S&P 500 datasets, which exhibit more unpredictable changes in volatility, the smoothness profile of ASV-DHS-N resembles that of MSSV2. The kurtosis remains substantially higher than that of SV and MSSV2, preserving the model’s ability to capture heavy-tailed behavior. The CP is also notably higher than that of SV, reflecting greater local adaptivity, though still lower than that of ASV-DHS.
@@
Figure 5: Comparison of the posterior means of the shrinkage parameter $\kappa_{t}:=1/(1+\exp\{v_{t}\})$, where $v_{t}=\log(\mathrm{Var}(\Delta h_{t}))$, and the log-variance parameter $h_{t}$, estimated using ASV-DHS-N on the S&P 500, electricity demand, and bike rental datasets. The shaded regions represent the 95% and centered 90% credible intervals for $\kappa_{t}$ and $h_{t}$, respectively. Vertical lines indicate dates where the expected $\kappa_{t}$ falls below 0.9.
@@
In Figure 5, we highlight time points where the posterior mean of $\kappa_{t}$ drops below 0.9, marking potential locations of abrupt volatility changes. For the S&P 500 index (Figure 5(a)), this includes notable episodes explored in earlier paragraph, such as the 2008 financial crisis, the 2011 European debt crisis, and the COVID-19 shock in early 2020. In Figure 5(b), electricity usage spikes in volatility around late spring and early fall, periods that typically mark transitions in heating or cooling demand, reflecting changing usage patterns driven by weather fluctuations. Figure 5(b) shows a similar pattern with volatility spikes in bike rental counts occurring near the start of spring such as late February and March and near the end-of-year holiday season in December and January.
@@ 6 Trend Filtering Jointly in Mean and Variance
The Bayesian Trend Filter with DSP (BTF-DSP) by Kowal et al. (2019) provides a smooth and locally adaptive estimate of the mean of a time series. As an extension, we propose the Bayesian Trend Filter with ASV (BTF-ASV), thereby allowing locally adaptive estimates of both the mean $\{\beta_{t}\}_{t=1}^{T}$, and the log variance process $\{h_{t}\}_{t=1}^{T}$ simultaneously:
@@
$y_{t}=\beta_{t}+\exp\{h_{t}/2\}\epsilon_{t}$ $\quad[\epsilon_{t}]\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1),$
$[\Delta^{k}\beta_{t}|\sigma^{2}_{\mu,t}]\sim N(0,\sigma_{\mu,t}^{2})$ $\quad[\Delta^{k}h_{t}|\sigma^{2}_{h,t}]\sim N(0,\sigma_{h,t}^{2}).$
@@
For contradiction, fix two distinct parameter vectors $\boldsymbol{\theta}:=(\boldsymbol{\beta},\boldsymbol{h},\boldsymbol{\lambda}_{\mu},\boldsymbol{\lambda}_{h})$ and $\boldsymbol{\theta}^{*}:=(\boldsymbol{\beta}^{*},\boldsymbol{h}^{*},\boldsymbol{\lambda}_{\mu}^{*},\boldsymbol{\lambda}_{h}^{*})$ such that $\boldsymbol{\theta}\neq\boldsymbol{\theta}^{*}$ and $f(y|\boldsymbol{\theta})=f(y|\boldsymbol{\theta}^{*})$. Indeed, $\mathrm{Var}(y|\boldsymbol{\theta})=\mathrm{Var}(y|\boldsymbol{\theta}^{*}).$ This covariance equality necessarily forces $\boldsymbol{\theta}=\boldsymbol{\theta}^{*}$, leading to contradiction. Detailed proof is shown in Appendix A. ∎
@@
Figure 6: Monthly mean total sunspot number, the average value of the daily total sunspot numbers over all days of each calendar month, from Jan. 1901 to Aug. 2024 is illustrated in 6(a). The Bayesian Trend Filter with Adaptive Stochastic Volatility (BTF-ASV) is used to estimate the mean trend $\beta_{t}$ and is illustrated in 6(b). The estimate of log-variance $h_{t}$ and its 90% credible intervals is shown in 6(c).
BTF-ASV is applied to the sunspot data to estimate both the time varying mean and the volatility. The data consist of monthly mean total sunspot numbers, defined as the average daily total over each calendar month, from January 1901 to August 2024, obtained from SILSO World Data Center (2025). The number of sunspots exhibits a well-known cyclic pattern associated with the 11-year solar cycle, which reflects periodic fluctuations in solar magnetic activity. This cyclical behavior is clearly visible in the raw observations shown in Figure 6(a). The BTF estimate of the mean trend, $\beta_{t}$, shown in Figure 6(b), provides a smooth reconstruction of this underlying cycle.
Interestingly, the estimated log-variance (Figure 6(c)) also follows a cyclical pattern that closely aligns with the mean trend. Periods of increased solar activity are associated with elevated volatility, suggesting that not only does solar activity rise and fall periodically, but the variability around these phases also intensifies. This highlights BTF-ASV’s ability to decompose a noisy data into structured patterns in both the mean and variance process, offering a richer understanding of solar dynamics.
@@ 7 Conclusion
We propose ASV, a flexible and locally adaptive model for capturing both abrupt and gradual changes in volatility. ASV extends RWSV by incorporating global-local shrinkage priors, specifically DSP by Kowal et al. (2019). We develop and study multiple variants of ASV, including models with the horseshoe (ASV-HS), dynamic horseshoe (ASV-DHS), and their nugget-augmented counterparts (ASV-HS-N and ASV-DHS-N).
Through extensive simulation studies, we show that ASV-HS-N and ASV-DHS-N consistently achieve low MAE, correct nominal coverage, and narrower credible intervals compared to existing volatility models. Although ASV variants without the nugget term perform similarly in terms of point estimation, they tend to exhibit under-coverage in stationary settings due to their lack of a noise-absorbing layer. Empirical applications demonstrate that ASV models effectively capture large structural changes in volatility while preserving smooth estimates in stable periods. Moreover, the shrinkage parameter $\kappa_{t}$ provides a useful diagnostic for further identifying the timing of structural shifts.
Finally, we integrate ASV into the error process of the Bayesian Trend Filter proposed in Kowal et al. (2019), resulting in BTF-ASV. Empirical studies reveal its ability to produce smooth estimates of both time-varying mean and time-varying variance. Future research directions could explore the applications of ASV and BTF-ASV in diverse fields such as finance, environmental science, and epidemiology, leveraging their adaptability in modeling time-series exhibiting complex patterns.
8 Acknowledgement
@@
9 Supplementary Material
Appendices:
A document including detailed proofs on the properties of Dynamic Shrinkage Processes (DSP) in section A. The exact prior distributions and detailed derivations of the full conditional distributions for Gibbs sampling in section B.
@@ A.1 Theorem 1
As mentioned, proofs directly follow from Kolmogorov’s three series theorem, which state the sufficient and necessary condition for almost sure convergence of infinite sum of random variables. Let $\phi\in(0,1)$. Due to symmetry of $sech$ function, we only need to prove the case when $10<\phi<$ 1, and the case where $-1<\phi<0$, directly follows. For the first condition, let $\epsilon>0$, and $h$ is some positive integer or 0.
[…]
$=\frac{4}{\pi}\text{arctan}(\text{exp}\{x/(2\phi^{h})\})\bigg{]}_{\epsilon}^{\infty}$
$=2(1-\frac{2}{\pi}\text{arctan}(\text{exp}\{\epsilon/(2\phi^{h})\})$
$=\frac{4}{\pi}\text{arccotan}(\text{exp}\{\epsilon/(2\phi^{h})\}$
$\leq\frac{4}{\pi}\text{arccotan}(\text{exp}\{h\})$ if $\textit{if }\epsilon\leq 2h\phi^{h}$
Since $10<\phi<$ 1, $2h\phi^{h}$ converges to 0. Also, $2h\phi^{h}\geq 0$ for $h\geq 0$. For any $\epsilon$, we have $h^{\prime}$ that satisfies $\epsilon\leq 2h\phi^{h}$ for $h^{\prime}\leq h$. Then, fix such $h^{\prime}$.
[…]
$\sum_{h=h^{\prime}}^{\infty}\frac{4}{\pi}\text{arccotan}(\text{exp}\{\epsilon/(2\phi^{h})\}\leq\sum_{h=h^{\prime}}^{\infty}\frac{4}{\pi}\text{arccotan}(\text{exp}\{h\})<\infty$
@@
$\sum_{h=0}^{\infty}\mathrm{P}(|z_{h}|\geq\epsilon)=\sum_{h=0}^{\infty}\frac{4}{\pi}\text{arccotan}(\text{exp}\{\epsilon/(2\phi^{h})\}<\infty$
On the side note, we can also easily see that when $|\phi|>1$, the sum diverges to $\infty$. By Borel-Cantelli lemma, we may conclude that $\sum_{h=0}^{\infty}z_{h}$ diverges almost surely.
Define $y_{h}=z_{h}1_{\{|z_{h}|\leq\epsilon\}}$. For the second and the third condition, we need to show that 1) $\sum_{h=0}^{\infty}\mathbb{E}(y_{h})$ converges, and 2) $\sum_{h=0}^{\infty}\mathrm{Var}(y_{h})$, converges. For the second condition, it suffices to show: E(∑_h=0^∞ y_h) = ∑_h=0^∞ E(y_h) ¡ ∞, which is satisfied by showing $\sum_{h=0}^{\infty}\mathbb{E}(|y_{h}|)<\infty$ (Fubini). Using Cauchy-Schwartz Inequality:
$\sum_{h=0}^{\infty}\mathbb{E}(|y_{h}|)$ $=\sum_{h=0}^{\infty}\mathbb{E}(|z_{h}1_{\{|z_{h}|\leq\epsilon\}}|)$
$\leq\sum_{h=0}^{\infty}\sqrt{\mathbb{E}(z_{h}^{2})\mathrm{P}(|z_{h}|\leq\epsilon)}$
$\leq\sum_{h=0}^{\infty}\sqrt{\mathbb{E}(z_{h}^{2})}=\sum_{h=0}^{\infty}\sqrt{\mathrm{Var}(z_{h})}=\pi\sum_{h=0}^{\infty}\phi^{h}=\frac{\pi}{(1-\phi)}<\infty.$
@@
$\mathrm{Var}(\sum_{h=0}^{\infty}y_{h})$ $=\mathbb{E}((\sum_{h=0}^{\infty}y_{h})^{2})+\mathbb{E}((\sum_{h=0}^{\infty}y_{h}))^{2}$
$=\mathbb{E}((\sum_{h=0}^{\infty}y_{h})^{2})+((\sum_{h=0}^{\infty}\mathbb{E}(y_{h}))^{2}$
$=\mathbb{E}((\sum_{h=0}^{\infty}y_{h})^{2}).$
@@
$\lim_{h\to\infty}\mathbb{E}((\sum_{n=0}^{n}y_{h})^{2})$ $=\lim_{n\to\infty}\mathbb{E}(\sum_{h=0}^{n}y_{h}^{2})$ $(independence)$
$=\lim_{n\to\infty}\sum_{h=0}^{n}\mathbb{E}(y_{h}^{2})=\sum_{h=0}^{\infty}\mathbb{E}(z_{h}^{2}1_{\{|z_{h}|\leq\epsilon\}})$
$\leq\sum_{h=0}^{\infty}\mathbb{E}(z_{h}^{2})=\frac{\pi^{2}}{(1-\phi^{2})}<\infty$
@@
$\mathrm{Var}(\sum_{h=0}^{\infty}y_{h})$ $=\sum_{h=0}^{\infty}\mathrm{Var}(y_{h})<\infty$

### v5

@@ 5 Empirical Study
We continue our analysis of the datasets introduced in Section 1 and estimate the log-variance using four methods: SV and MSSV2, which serve as established baselines, and ASV-DHS and ASV-DHS-N, our proposed models. For the ASV models, we set $k=1$ to reflect the presence of abrupt changes observed in Figure 1. In the S&P 500 dataset (Figures 4(a), 4(b), 4(c) and 4(d)), all models detect major volatility spikes in 2009 and 2020, corresponding to the global financial crisis and the COVID-19 pandemic. Moderate increases are also observed around 2001 (9/11), 2011 (European debt crisis), and 2018 (U.S. government shutdowns). ASV-DHS and ASV-DHS-N provide smoother and more stable estimates compared to SV and MSSV2, particularly during low-volatility periods. For both the electricity (Figures 4(e), 4(f), 4(g) and 4(h)) and bike rental datasets (Figures 4(i), 4(j), 4(k) and 4(l)), all methods capture strong seasonal patterns in volatility, with elevated uncertainty coinciding with periods of peak activity. In the electricity dataset, this corresponds to high-demand months, while in the bike rental dataset, volatility increases during the summer.
@@
$\mathrm{LL}$ 2.708 2.710 2.648 2.564
Table 2: Empirical results: (1) the mean and (2) kurtosis of the estimated log-variance increments $|\Delta\hat{h}_{t}|$ and (3) the total number of flagged change points (CP) and (4) Logarithmic Loss ($L$) on short term forecasts (6 step-ahead). CP are computed as $\sum_{t=2}^{T}\mathds{1}\{|\Delta\hat{h}_{t}|>\delta\}$, where $\delta:=5\times\texttt{sd}(\Delta\hat{h}_{t})$. $L$ is computed as $\frac{1}{6}\sum_{t=1}^{6}\left|\log(y_{T+t}^{2})-\hat{h}_{T+t}\right|$, where $T$ denotes the length of the training sample and $\hat{h}_{t}$ is the estimated log-variance. Each model is first estimated using the first 100 observations, then re-estimated using the first 200, 300, and so on, with a 6-step-ahead forecast generated at each stage until the end of the sample.
Table 2 presents summary metrics of the estimated log-variance increments $|\Delta\hat{h}_{t}|$ across the three empirical datasets. These metrics, mean, kurtosis, the number of flagged change points (CP), and the logarithmic loss (LL) computed from short-term (6-step-ahead) forecasts. The first three metrics offer insight into the smoothness, tail behavior, and local adaptivity, while LL offers a rough indication of short-term predictive accuracy (pagan).
ASV-DHS consistently produces the lowest mean, indicating that, on average, it exhibits the smallest changes in the estimated volatility compared to other models. At the same time, it shows extremely high kurtosis and the highest number of flagged change points, suggesting heavy-tailed behavior in the increments and strong local adaptivity to sudden structural shifts. These properties highlight the model’s tendency to promote overall smoothness while remaining responsive to large, isolated changes.
The smoothness and local adaptability of ASV-DHS-N vary depending on the underlying structure of the data. In the electricity dataset, where volatility is characterized by extended periods of small changes punctuated by occasional abrupt shifts, the mean in Table 2 under ASV-DHS-N is similar to that of ASV-DHS, indicating comparable overall variation. In contrast, for the bike rental and S&P 500 datasets, which exhibit more unpredictable changes in volatility, the mean and CP patterns resemble those of MSSV2, suggesting behavior similar to a smoother regime-switching model. The kurtosis remains substantially higher than that of SV and MSSV2, preserving the model’s ability to capture heavy-tailed behavior.
The LL results indicate that the proposed models, ASV-DHS and ASV-DHS-N, despite not being developed as forecasting models, perform comparably to benchmark approaches overall, with slightly lower LL for the electricity and bike rental datasets. The notably lower LL values for the electricity and bike rental data are consistent with their clear nonstationary volatility patterns, characterized by abrupt and irregular changes. In contrast, the smaller differences among models for the S&P 500 data suggest comparatively more stable volatility dynamics over the sample period.
@@
Figure 5: Comparison of the posterior means of the shrinkage parameter, $\kappa_{t}:=1/(1+\exp{v_{t}})$, where $v_{t}=\log(\mathrm{Var}(\Delta h_{t}))$, and the log-variance parameter, $h_{t}$, estimated using ASV-DHS-N for the S&P 500, electricity demand, and bike rental datasets. The shaded regions represent the 95% and centered 90% credible intervals for $\kappa_{t}$ and $h_{t}$, respectively. Vertical lines indicate the dates at which the expected $\kappa_{t}$ falls below 0.9.
@@
In Figure 5, we highlight the time points in which the posterior mean of $\kappa_{t}$ falls below 0.9, marking potential periods of abrupt volatility changes. For the S&P 500 index (Figure 5(a)), this includes notable episodes explored earlier, such as the 2008 financial crisis, the 2011 European debt crisis, and the COVID-19 shock in early 2020. In Figure 5(b), volatility in electricity usage spikes around late spring and early fall, periods that typically mark transitions in heating or cooling demand, reflecting changing usage patterns driven by weather fluctuations. Figure 5(c) shows a similar pattern with volatility spikes in bike rental counts occurring near the start of spring, such as late February and March, and near the end-of-year holiday season in December and January.
@@ 6 Trend Filtering Jointly in Mean and Variance
The Bayesian Trend Filter with DSP (BTF-DSP) by dsp provides a smooth and locally adaptive estimate of the mean of a time series. As an extension, we propose the Bayesian Trend Filter with ASV (BTF-ASV), thereby allowing locally adaptive estimates of both the mean $\{\beta_{t}\}_{t=1}^{T}$, and the log-variance process $\{h_{t}\}_{t=1}^{T}$ simultaneously:
@@
$y_{t}=\beta_{t}+\exp\{h_{t}/2\}\epsilon_{t}$ $\qquad[\epsilon_{t}]\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1),$
$[\Delta^{k}\beta_{t}|\sigma^{2}_{\beta,t}]\sim N(0,\sigma_{\beta,t}^{2})$ $\qquad[\Delta^{k}h_{t}|\sigma^{2}_{h,t}]\sim N(0,\sigma_{h,t}^{2}).$
@@
Let $\boldsymbol{\beta}:=\{\beta_{1},\ldots,\beta_{T}\}$ and $\boldsymbol{h}:=\{h_{1},\ldots,h_{T}\}$ denote the latent mean and log-volatility processes, respectively. For notational convenience, define $\boldsymbol{\lambda}_{\beta}:=\{\sigma_{\beta,1}^{2},\ldots,\sigma_{\beta,T}^{2}\}$ and $\boldsymbol{\lambda}_{h}:=\{\sigma_{h,1}^{2},\ldots,\sigma_{h,T}^{2}\}$. For contradiction, fix two distinct parameter vectors $\boldsymbol{\theta}:=(\boldsymbol{\beta},\boldsymbol{h},\boldsymbol{\lambda}_{\beta},\boldsymbol{\lambda}_{h})$ and $\boldsymbol{\theta}^{*}:=(\boldsymbol{\beta}^{*},\boldsymbol{h}^{*},\boldsymbol{\lambda}_{\beta}^{*},\boldsymbol{\lambda}_{h}^{*})$ such that $\boldsymbol{\theta}\neq\boldsymbol{\theta}^{*}$ and $f(y|\boldsymbol{\theta})=f(y|\boldsymbol{\theta}^{*})$. Indeed, $\mathrm{Var}(y|\boldsymbol{\theta})=\mathrm{Var}(y|\boldsymbol{\theta}^{*}).$ This covariance equality necessarily forces $\boldsymbol{\theta}=\boldsymbol{\theta}^{*}$, leading to contradiction. Detailed proof is shown in Appendix A. ∎
While Theorem 4 establishes the theoretical identifiability of the BTF-ASV model, practical identifiability, the extent to which the available data are informative enough to separate the first-order effect from the second-order effects, may still be limited under severe model misspecification or high noise in the analyzed data. Practical identifiability in Bayesian models is an active area of research (pident1; pident3; pident2). Extending such ideas to time-series models may be an interesting direction for future research.
@@
Figure 6: Monthly mean total sunspot numbers, defined as the average of daily total sunspot counts over each calendar month from January 1901 to August 2024, are shown in Figure 6(a). The Bayesian Trend Filter with Adaptive Stochastic Volatility (BTF-ASV) is used to estimate the mean trend, $\beta_{t}$, as shown in Figure 6(b). The estimated log-variance, $h_{t}$, along with its 90% credible intervals, is presented in Figure 6(c).
BTF-ASV is applied to the sunspot data to estimate both the time-varying mean and the volatility. The data consist of monthly mean total sunspot numbers, defined as the average daily total over each calendar month, from January 1901 to August 2024, obtained from sunspot. The number of sunspots exhibits a well-known cyclic pattern associated with the 11-year solar cycle, which reflects periodic fluctuations in solar magnetic activity. This cyclical behavior is clearly visible in the raw observations shown in Figure 6(a). The BTF-ASV estimate of the mean trend, $\beta_{t}$, shown in Figure 6(b), provides a smooth reconstruction of this underlying cycle.
Interestingly, the estimated log-variance (Figure 6(c)) also follows a cyclical pattern that closely aligns with the mean trend. Periods of increased solar activity are associated with elevated volatility, suggesting that not only does solar activity rise and fall periodically, but the variability around these phases also intensifies. This highlights BTF-ASV’s ability to decompose noisy data into structured patterns in both the mean and variance processes, offering a richer understanding of solar dynamics.
@@ 7 Conclusion
We propose ASV, a flexible and locally adaptive model for capturing both abrupt and gradual changes in volatility. ASV extends RWSV by incorporating a global-local shrinkage prior, specifically the DSP by dsp. We develop and study multiple variants of ASV, including models with the horseshoe (ASV-HS), dynamic horseshoe (ASV-DHS), and their nugget-augmented counterparts (ASV-HS-N and ASV-DHS-N).
Through extensive simulation studies, we show that ASV-HS-N and ASV-DHS-N consistently achieve low MAE, correct nominal coverage, and narrower credible intervals compared to existing models. Empirical applications demonstrate that ASV models effectively capture large structural changes in volatility while preserving smooth estimates in stable periods. Moreover, the shrinkage parameter $\kappa_{t}$ provides a useful diagnostic for further identifying the timing of structural shifts.
Finally, we integrate ASV into the error process of the Bayesian Trend Filter proposed in dsp, resulting in the BTF-ASV model. Empirical studies reveal its ability to produce smooth and locally adaptive estimates of both the time-varying mean and variance. Future research directions could explore the applications of ASV and BTF-ASV in diverse fields such as finance, environmental science, and epidemiology, leveraging their capacity to produce adaptive estimates for time-series that exhibit abrupt changes in both the mean and variance.
8 Acknowledgment
@@
9 Disclosure Statement
The authors report there are no competing interests to declare.
@@ A.1 Theorem 1
This proof directly follows from Kolmogorov’s three series theorem, which states the sufficient and necessary condition for almost sure convergence of infinite sum of random variables. Let $\phi\in(0,1)$. Due to symmetry of sech function, we only need to prove the case when $0<\phi<1$, and the case where $1<\phi<0$, directly follows. For the first condition, let $\epsilon>0$, and $\ell$ is some positive integer or 0.
$\mathrm{P}(|z_{\ell}|\geq\epsilon)$ $=2\int_{\epsilon}^{\infty}\frac{\text{exp}\{x/(2\phi^{\ell})\}}{\pi\phi^{\ell}(1+\text{exp}\{x/\phi^{\ell}\}}dx$
$=\frac{2}{\pi\phi^{\ell}}\int_{\epsilon}^{\infty}\frac{\text{exp}\{x/(2\phi^{\ell})\}}{1+\text{exp}\{x/\phi^{\ell}\}}dx$ $u=\text{exp}\{x/(2\phi^{\ell})\}$
$=\frac{4}{\pi}\int_{\epsilon}^{\infty}\frac{1}{1+u^{2}}du$ $2\phi^{\ell}du=\text{exp}\{x/(2\phi^{\ell})\}dx$
$=\frac{4}{\pi}\text{arctan}(\text{exp}\{x/(2\phi^{\ell})\})\bigg]_{\epsilon}^{\infty}$
$=2(1-\frac{2}{\pi}\text{arctan}(\text{exp}\{\epsilon/(2\phi^{\ell})\})$
$=\frac{4}{\pi}\text{arccotan}(\text{exp}\{\epsilon/(2\phi^{\ell})\}$
$\leq\frac{4}{\pi}\text{arccotan}(\text{exp}\{\ell\})$ $\textit{if }\epsilon\leq 2\ell\phi^{\ell}$
Since $0<\phi<1$, $2\ell\phi^{\ell}$ converges to 0. Also, $2\ell\phi^{l}\geq 0$ for $\ell\geq 0$. For any $\epsilon$, we have $\ell^{\prime}$ that satisfies $\epsilon\leq 2\ell\phi^{\ell}$ for $\ell^{\prime}\leq\ell$. Then, fix such $\ell^{\prime}$.
$\int_{\ell^{\prime}}^{\infty}\frac{4}{\pi}\text{arccotan}(\text{exp}\{\ell\})dh<\infty$
$\sum_{\ell=\ell^{\prime}}^{\infty}\frac{4}{\pi}\text{arccotan}(\text{exp}\{\epsilon/(2\phi^{\ell})\}\leq\sum_{\ell=\ell^{\prime}}^{\infty}\frac{4}{\pi}\text{arccotan}(\text{exp}\{\ell\})<\infty$
@@
$\sum_{\ell=0}^{\infty}\mathrm{P}(|z_{\ell}|\geq\epsilon)=\sum_{h=0}^{\infty}\frac{4}{\pi}\text{arccotan}(\text{exp}\{\epsilon/(2\phi^{\ell})\}<\infty$
On the side note, we can also easily see that when $|\phi|>1$, the sum diverges to $\infty$. By Borel-Cantelli lemma, we may conclude that $\sum_{\ell=0}^{\infty}z_{\ell}$ diverges almost surely.
Define $y_{\ell}=z_{\ell}1_{\{|z_{\ell}|\leq\epsilon\}}$. For the second and the third condition, we need to show that 1) $\sum_{\ell=0}^{\infty}\mathbb{E}(y_{\ell})$ converges, and 2) $\sum_{\ell=0}^{\infty}\mathrm{Var}(y_{\ell})$, converges. For the second condition, it suffices to show: E(∑_ℓ=0^∞ y_ℓ) = ∑_ℓ=0^∞ E(y_ℓ) ¡ ∞, which is satisfied by showing $\sum_{\ell=0}^{\infty}\mathbb{E}(|y_{\ell}|)<\infty$ (Fubini). Using Cauchy-Schwartz Inequality:
$\sum_{\ell=0}^{\infty}\mathbb{E}(|y_{\ell}|)$ $=\sum_{\ell=0}^{\infty}\mathbb{E}(|z_{\ell}1_{\{|z_{\ell}|\leq\epsilon\}}|)$
$\leq\sum_{\ell=0}^{\infty}\sqrt{\mathbb{E}(z_{\ell}^{2})\mathrm{P}(|z_{\ell}|\leq\epsilon)}$
$\leq\sum_{\ell=0}^{\infty}\sqrt{\mathbb{E}(z_{\ell}^{2})}=\sum_{\ell=0}^{\infty}\sqrt{\mathrm{Var}(z_{\ell})}=\pi\sum_{\ell=0}^{\infty}\phi^{\ell}=\frac{\pi}{(1-\phi)}<\infty.$
@@
$\mathrm{Var}\bigg(\sum_{\ell=0}^{\infty}y_{\ell}\bigg)$ $=\mathbb{E}\bigg(\bigg(\sum_{\ell=0}^{\infty}y_{\ell}\bigg)^{2}\bigg)+\bigg(\mathbb{E}\bigg(\sum_{\ell=0}^{\infty}y_{\ell}\bigg)\bigg)^{2}$
$=\mathbb{E}\bigg(\bigg(\sum_{\ell=0}^{\infty}y_{\ell}\bigg)^{2}\bigg)+\bigg(\sum_{\ell=0}^{\infty}\mathbb{E}(y_{\ell})\bigg)^{2}.$
@@
$\lim_{n\to\infty}\mathbb{E}\bigg(\bigg(\sum_{\ell=0}^{n}y_{\ell}\bigg)^{2}\bigg)$ $=\lim_{n\to\infty}\mathbb{E}\bigg(\sum_{\ell=0}^{n}y_{\ell}^{2}\bigg)$ $(independence)$
$=\lim_{n\to\infty}\sum_{\ell=0}^{n}\mathbb{E}(y_{\ell}^{2})=\sum_{\ell=0}^{\infty}\mathbb{E}(z_{\ell}^{2}1_{\{|z_{\ell}|\leq\epsilon\}})$
$\leq\sum_{\ell=0}^{\infty}\mathbb{E}(z_{\ell}^{2})=\frac{\pi^{2}}{(1-\phi^{2})}<\infty$
@@
$\mathrm{Var}(\sum_{\ell=0}^{\infty}y_{\ell})$ $=\sum_{\ell=0}^{\infty}\mathrm{Var}(y_{\ell})<\infty$